# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ Appendix A Proofs of Lemmas
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=\sinh h_{i}$, then $\cosh q_{k}=-\sinh h_{i}\cdot\sinh d_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, then $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$. If $c^{\prime}_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=+\sinh(-q_{j})=-\sinh q_{j}$. If $c^{\prime}_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=-\sinh(+q_{j})=-\sinh q_{j}$. Thus $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$. Then $\cosh h_{i}=-\sinh q_{j}\cdot\sinh\theta_{ik}$. By Lemma 2.6, it is easy to check that $\theta_{ik}<0$. Hence $q_{j}>0$. Then $d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot})=q_{j}$, which implies that $c^{\prime}_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$. If $c^{\prime}_{ij}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=+\sinh(-q_{k})=-\sinh q_{k}$. If $c^{\prime}_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=-\sinh(+q_{k})=-\sinh q_{k}$. Thus $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$. Then $\cosh h_{i}=-\sinh q_{k}\cdot\sinh\theta_{ij}$. Similarly, $\theta_{ij}<0$ by Lemma 2.6 and hence $q_{k}>0$. Then $d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot})=q_{k}$, which implies that $c^{\prime}_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$ and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$. Then $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$ and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, then $\cosh h_{k}=\sinh q_{i}\cdot\sinh\theta_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 $\mathrm{(iii)}$, $v_{i}\ast c_{ijk}=\cosh d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot})=\cosh q_{i}$. By Proposition 2.2 $\mathrm{(ii)}$, $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$. If $c_{ik}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh h_{j}$; If $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh(-h_{j})=+\sinh h_{j}$. Thus $c_{ik}\ast c_{ijk}=\sinh h_{j}$. Then $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$. By Lemma 2.5, it is easy to check that $d_{ik}<0$. Then $h_{j}<0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=-h_{j}$, which implies that $c_{ik}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})$. Similarly, $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$. Then $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{jk}=-\sinh d_{jk}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, then $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$ and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, then $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$ and $c_{ik}\ast c_{ijk}=\sinh h_{j}$, then $\cosh q_{k}=-\sinh h_{j}\cdot\sinh d_{ki}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$ and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, then $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$ and $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$. If $c^{\prime}_{jk}$ and $c_{ijk}$ are on same sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=+\sinh q_{i}$; If $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$. Thus $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$. Then $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$. By Lemma 2.6, it is easy to check that $\theta_{jk}<0$. Then $q_{i}<0$ and hence $d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot})=-q_{i}$, which implies that $c^{\prime}_{jk}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$ and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, then $\cosh h_{j}=-\sinh q_{k}\cdot\sinh\theta_{ji}$.

### v3

@@ Appendix A Proofs of Lemmas
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})=\cosh q_{i}$. Since $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{i}=\sinh h_{k}\cdot\sinh d_{ij}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. By Proposition 2.2 (iii), $v_{j}\ast c_{ijk}=-\cosh(d_{\mathbb{H}}(v_{j}^{\bot},c_{ijk}^{\bot}))=-\cosh q_{j}$. Since $c_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, it follows that $c_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{jk},c_{ijk}^{\bot}))=-\sinh(-h_{i})=+\sinh h_{i}$. Due to $v_{j}\ast c_{jk}=-\sinh d_{jk}$, we conclude that $\cosh q_{j}=-\sinh h_{i}\cdot\sinh d_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$, and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$, and $c_{jk}\ast c_{ijk}=\sinh h_{i}$, it follows that $\cosh q_{k}=-\sinh h_{i}\cdot\sinh d_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$, and $c_{ik}\ast c_{ijk}=-\sinh h_{j}$, it follows that $\cosh q_{k}=\sinh h_{j}\cdot\sinh d_{ki}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$. If $c^{\prime}_{ik}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=+\sinh(-q_{j})=-\sinh q_{j}$. If $c^{\prime}_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot}))=-\sinh(+q_{j})=-\sinh q_{j}$. In either case, $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$. Consequently, $\cosh h_{i}=-\sinh q_{j}\cdot\sinh\theta_{ik}$. By Lemma 2.6, it is easy to check that $\theta_{ik}<0$. It follows that $q_{j}>0$ and hence $d_{\mathbb{H}}(c^{\prime}_{ik},c_{ijk}^{\bot})=q_{j}$. This implies that $c^{\prime}_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Similarly, $v^{\prime}_{i}\ast c_{ijk}=\cosh h_{i}$ and $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$. If $c^{\prime}_{ij}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=+\sinh(-q_{k})=-\sinh q_{k}$. If $c^{\prime}_{ij}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{ij}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot}))=-\sinh(+q_{k})=-\sinh q_{k}$. In either case, $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$. Consequently, $\cosh h_{i}=-\sinh q_{k}\cdot\sinh\theta_{ij}$. Similarly, $\theta_{ij}<0$ by Lemma 2.6 and hence $q_{k}>0$. Then $d_{\mathbb{H}}(c^{\prime}_{ij},c_{ijk}^{\bot})=q_{k}$, which implies that $c^{\prime}_{ij}$ and $c_{ijk}$ are on opposite sides of the hyperplane $c_{ijk}^{\bot}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$, and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, it follows that $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{j}\ast c_{ijk}=-\cosh h_{j}$, $v^{\prime}_{j}\ast c^{\prime}_{ij}=-\sinh\theta_{ji}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{j}=\sinh q_{k}\cdot\sinh\theta_{ji}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{jk})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{jk}=-\sinh\theta_{kj}$, and $c^{\prime}_{jk}\ast c_{ijk}=-\sinh q_{i}$, it follows that $\cosh h_{k}=\sinh q_{i}\cdot\sinh\theta_{kj}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{k}$, we have $-v^{\prime}_{k}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{k}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{k}\ast c_{ijk}=-\cosh h_{k}$, $v^{\prime}_{k}\ast c^{\prime}_{ik}=-\sinh\theta_{ki}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{k}=\sinh q_{j}\cdot\sinh\theta_{ki}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{i}\ast c_{ik})$. By Proposition 2.2 (iii), $v_{i}\ast c_{ijk}=\cosh d_{\mathbb{H}}(v_{i}^{\bot},c_{ijk}^{\bot})=\cosh q_{i}$. By Proposition 2.2 (ii), $v_{i}\ast c_{ik}=+\sinh d_{\mathbb{H}}(v_{i}^{\bot},c_{ik})=+\sinh(-d_{ik})=-\sinh d_{ik}$. If $c_{ik}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=+\sinh h_{j}$; if $c_{ik}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c_{ik}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot}))=-\sinh(-h_{j})=+\sinh h_{j}$. In either case, $c_{ik}\ast c_{ijk}=\sinh h_{j}$. Consequently, $\cosh q_{i}=\sinh h_{j}\cdot\sinh d_{ik}$. By Lemma 2.5, it is easy to check that $d_{ik}<0$. It follows that $h_{j}<0$ and hence $d_{\mathbb{H}}(c_{ik},c_{ijk}^{\bot})=-h_{j}$. This implies that $c_{ik}$ and $c_{ijk}$ lie on opposite sides of $c_{ijk}^{\bot}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{i}$, we have $-v_{i}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{i}\ast c_{ij})$. Since $v_{i}\ast c_{ij}=-\sinh d_{ij}$ and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{i}=-\sinh h_{k}\cdot\sinh d_{ij}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{j}\ast c_{jk})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{jk}=-\sinh d_{jk}$, and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, it follows that $\cosh q_{j}=\sinh h_{i}\cdot\sinh d_{jk}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ij}v_{j}$, we have $-v_{j}\ast c_{ijk}=(c_{ij}\ast c_{ijk})(v_{j}\ast c_{ij})$. Since $v_{j}\ast c_{ijk}=-\cosh q_{j}$, $v_{j}\ast c_{ij}=-\sinh d_{ji}$, and $c_{ij}\ast c_{ijk}=-\sinh h_{k}$, it follows that $\cosh q_{j}=\sinh h_{k}\cdot\sinh d_{ji}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{jk}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{jk}\ast c_{ijk})(v_{k}\ast c_{jk})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{jk}=-\sinh d_{kj}$, and $c_{jk}\ast c_{ijk}=-\sinh h_{i}$, it follows that $\cosh q_{k}=\sinh h_{i}\cdot\sinh d_{kj}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c_{ik}v_{k}$, we have $-v_{k}\ast c_{ijk}=(c_{ik}\ast c_{ijk})(v_{k}\ast c_{ik})$. Since $v_{k}\ast c_{ijk}=-\cosh q_{k}$, $v_{k}\ast c_{ik}=-\sinh d_{ki}$, and $c_{ik}\ast c_{ijk}=\sinh h_{j}$, it follows that $\cosh q_{k}=-\sinh h_{j}\cdot\sinh d_{ki}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ik}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ik}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ik})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ik}=-\sinh\theta_{ik}$, and $c^{\prime}_{ik}\ast c_{ijk}=-\sinh q_{j}$, it follows that $\cosh h_{i}=\sinh q_{j}\cdot\sinh\theta_{ik}$.
@@
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{ij}v^{\prime}_{i}$, we have $-v^{\prime}_{i}\ast c_{ijk}=(c^{\prime}_{ij}\ast c_{ijk})(v^{\prime}_{i}\ast c^{\prime}_{ij})$. Since $v^{\prime}_{i}\ast c_{ijk}=-\cosh h_{i}$, $v^{\prime}_{i}\ast c^{\prime}_{ij}=-\sinh\theta_{ij}$, and $c^{\prime}_{ij}\ast c_{ijk}=-\sinh q_{k}$, it follows that $\cosh h_{i}=\sinh q_{k}\cdot\sinh\theta_{ij}$.
In the generalized right-angled hyperbolic triangle $c_{ijk}c^{\prime}_{jk}v^{\prime}_{j}$, we have $-v^{\prime}_{j}\ast c_{ijk}=(c^{\prime}_{jk}\ast c_{ijk})(v^{\prime}_{j}\ast c^{\prime}_{jk})$. Moreover, $v^{\prime}_{j}\ast c_{ijk}=\cosh h_{j}$ and $v^{\prime}_{j}\ast c^{\prime}_{jk}=-\sinh\theta_{jk}$. If $c^{\prime}_{jk}$ and $c_{ijk}$ lie on the same side of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=+\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=+\sinh q_{i}$; if $c^{\prime}_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$, then $c^{\prime}_{jk}\ast c_{ijk}=-\sinh(d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot}))=-\sinh(-q_{i})=+\sinh q_{i}$. In either case, $c^{\prime}_{jk}\ast c_{ijk}=\sinh q_{i}$. Consequently, $\cosh h_{j}=\sinh q_{i}\cdot\sinh\theta_{jk}$. By Lemma 2.6, it is easy to check that $\theta_{jk}<0$. It follows that $q_{i}<0$ and hence $d_{\mathbb{H}}(c^{\prime}_{jk},c_{ijk}^{\bot})=-q_{i}$. This implies that $c^{\prime}_{jk}$ and $c_{ijk}$ lie on opposite sides of the hyperplane $c_{ijk}^{\bot}$.